OpenStudy (anonymous):

(square root of 2( cos 10+ i sin 10 ))^6 in standard form

OpenStudy (xishem):

Is it:\[\left(\sqrt{2(\cos(10)+i*\sin(10))} \right)^6\]?

OpenStudy (anonymous):

yes

OpenStudy (xishem):

OK, so first, the sum of trig terms can be simplified into an exponential using Euler's relation. Do you know it?